# RetrySQL: text-to-SQL training with retry data for self-correcting query generation

### v1

@@ Abstract.
The text-to-SQL task is an active challenge in Natural Language Processing. Many existing solutions focus on using black-box language models extended with specialized components within customized end-to-end text-to-SQL pipelines. While these solutions use both closed-source proprietary language models and coding-oriented open-source models, there is a lack of research regarding SQL-specific generative models. At the same time, recent advancements in self-correcting generation strategies show promise for improving the capabilities of existing architectures. The application of these concepts to the text-to-SQL task remains unexplored. In this paper, we introduce RetrySQL, a new approach to training text-to-SQL generation models. We prepare reasoning steps for reference SQL queries and then corrupt them to create retry data that contains both incorrect and corrected steps, divided with a special token. We continuously pre-train an open-source coding model with this data and demonstrate that retry steps yield an improvement of up to 4 percentage points in both overall and challenging execution accuracy metrics, compared to pre-training without retry data. Additionally, we confirm that supervised fine-tuning with LoRA is ineffective for learning from retry data and that full-parameter pre-training is a necessary requirement for that task. We showcase that the self-correcting behavior is learned by the model and the increase in downstream accuracy metrics is a result of this additional skill. Finally, we incorporate RetrySQL-trained models into the full text-to-SQL pipeline and showcase that they are competitive in terms of execution accuracy with proprietary models that contain orders of magnitude more parameters. RetrySQL demonstrates that self-correction can be learned in the text-to-SQL task and provides a novel way of improving generation accuracy for SQL-oriented language models.
text-to-SQL, SQL, retry data, self-correction, continued pre-training, large language model
ccs: Computing methodologies Natural language processing
1. Introduction
@@
The task of translating natural language questions to SQL queries is a major challenge for machine learning models. The complexity stems from the need of relating often ambiguous user input to abstract entities, relations and values that are present in relational databases (Li et al., 2023). Even prominent Large Language Models (LLMs), such as GPT-4o (Team, 2024b) or Gemini 1.5 (Team, 2024a), struggle with approaching human performance in leading text-to-SQL benchmarks: BIRD (Li et al., 2023) and SPIDER 2.0 (Lei et al., 2024). The same is true for models tuned specifically for coding tasks (Li et al., 2024; Chen et al., 2024; Talaei et al., 2024). Thus, there exists a need for more advanced solutions that can bridge that gap and provide reliable SQL queries even in difficult real-world scenarios.
@@
The text-to-SQL task can be divided into three main steps (Maamari et al., 2024): retrieval, generation and correction. Many existing approaches try to tackle these steps at the same time, in a single end-to-end pipeline (Maamari et al., 2024; Talaei et al., 2024; Pourreza and Rafiei, 2023, 2024). In this work, we focus only on the generation step and show how it can be improved with a novel approach to model pre-training.
@@
Figure 1. RetrySQL overview. (a) Reasoning step generation. For each SQL query in the training dataset, we generate a series of reasoning steps using GPT-4o. (b) Preparation of retry data. For each set of reasoning steps, we apply random perturbations, treated as errors, by replacing some steps with different ones. We follow these errors with special [BACK] tokens and amend them with correct steps. (c) We take an open-source LLM and continue its pre-training with training examples that contain retry data injected into reasoning steps. The resulting RetrySQL-trained model learns the ability to self-correct, which improves its capabilities in generating correct SQL queries from natural language questions.
@@
Specifically, we teach the model to self-correct during the generation itself. While previous work did use self-correction in the sense of post-processing, applied in the correction step (Pourreza and Rafiei, 2023), we enforce the self-correcting behavior at an earlier point. This sort of active knowledge-based self-correction is an ongoing research area when it comes to LLMs (Ye et al., 2024a, b). While recent work in slow thinking reasoning systems, such as DeepSeek-R1 (Team, 2025), shows that self-correction can be learned in a reinforcement learning setup, other lines of research suggest that is is possible to obtain the self-correction ability with specific data augmentations and a standard auto-regressive pre-training objective (Ye et al., 2024a, b). It has been shown that augmenting training data for grade-school math solution generation with so-called retry data leads to increased generation accuracy (Ye et al., 2024b). The applicability of this approach to other tasks and models has not been explored as of yet.
@@
We introduce RetrySQL, a novel text-to-SQL generation module training paradigm that incorporates retry data in the training process and teaches the resulting model to self-correct. RetrySQL first augments the training data with reasoning steps that explain the sequence of operations required for obtaining the solution SQL query (Fig. 1a). Then, retry data is generated by corrupting the order of these reasoning steps (Fig. 1b). The retry data is incorporated into the training examples and we perform continued pre-training of an open-source coding-oriented LLM, which results in a RetrySQL-trained model that is capable of self-correction (Fig. 1c). We present a set of experiments across multiple strategies of generating reasoning corruptions. We demonstrate that using retry data yields superior generation results when compared to training data with error-free reasoning steps.
@@
To demonstrate that the self-correcting behavior is indeed learned by the model, we provide an analysis of output token confidence. We show that the max softmax score is on average lower for tokens before correction than for those after. Similarly, incorrect tokens display higher variance of softmax scores across beam search passes than the corrected ones. This illustrates that the model becomes uncertain as it makes a mistake and then self-corrects itself with higher confidence.
The standard approach to adapting open-source LLMs to specific tasks is supervised fine-tuning (SFT) with Low-Rank Adaptation (LoRA) (Hu et al., 2022), which is a byproduct of large model sizes and limited compute resources for most machine learning practitioners. We show that SFT with LoRA is insufficient for learning the self-correction ability from retry data and that full- parameter continued pre-training is a necessary requirement for that task.
@@
Our results corroborate the recent findings regarding the self-correction ability in language models (Ye et al., 2024b), demonstrating that the improvement in generation accuracy coming from the inclusion of retry data in pre-training is a universal law. It is applicable not only to the grade school math reasoning problem and GPT-2, but also to the text-to-SQL domain and larger, more modern Transformer-based decoder-only models. These findings suggest that retry data could be adapted to even more domains, especially if reasoning steps can be added to the training examples.
@@
While the main focus of our work is on the SQL generation step in isolation, we also showcase that relatively small, 1.5B-parameter open-source coding models trained with our RetrySQL paradigm are competitive with much larger closed-source proprietary LLMs when used as a part of the full text-to-SQL pipeline. We share all code for retry data generation, as well as model training and evaluation [^1] https://github.com/allegro/RetrySQL.
@@
  • We show that using retry data in pre-training is beneficial to the generation process, as indicated by the Execution Accuracy metric calculated for the BIRD benchmark dataset.
@@
  • We show that full-parameter continued pre-training is necessary for retry data in the text-to-SQL task, as opposed to the typical LoRA-based supervised fine-tuning.
@@
  • We illustrate that within a simple end-to-end text-to-SQL pipeline, RetrySQL-trained 1.5B-parameter open-source coding models are competitive with proprietary models such as GPT-4o-mini and GPT-4o.
2. Related work
@@
Early text-to-SQL methods relied on sequence-to-sequence frameworks, using models like Graph Neural Networks, Recurrent Neural Networks, and pre-trained Transformers for encoding queries and schemas (Cai et al., 2021; Cao et al., 2021; Hwang et al., 2019), while employing slot-filling or auto-regressive decoding to generate SQL queries (Choi et al., 2020; Wang et al., 2020). Recently, the field has shifted with the emergence of LLMs, which are currently leading in the most popular benchmarks (Shi et al., 2024). While initial efforts were focused on optimizing prompt designs that leveraged in-context learning (Nan et al., 2023; Gao et al., 2023) and multi-stage prompting (Pourreza and Rafiei, 2023), the current state-of-the-art is represented by LLM-based pipelines. These latest approaches integrate LLMs in more complex sequences of processing stages, with separate components for schema linking, self-correction, self-debugging, and self-consistency (Talaei et al., 2024; Pourreza et al., 2024; Lee et al., 2024; Maamari et al., 2024; Sun et al., 2024).
@@
Compared to closed-source model prompting approaches, open-source model fine-tuning for the text-to-SQL task remains relatively unexplored (Shi et al., 2024). Many of the existing works favor parameter-efficient fine-tuning (PEFT) over full-parameter fine-tuning due to the former’s superior training efficiency and lower training costs (Shi et al., 2024; Chen et al., 2024; Zhang et al., 2024a, b). While the majority of practitioners choose to use powerful general-purpose LLMs as their base models (Pourreza and Rafiei, 2024; Xie et al., 2024; Shi et al., 2024), promising results have also been shown by adapting coding LLMs to the text-to-SQL domain (Li et al., 2024; Chen et al., 2024; Talaei et al., 2024), demonstrating that starting from a model already heavily pre-trained on coding tasks, with SQL-related training data, leads to higher performance in benchmark evaluations.
The bulk of text-to-SQL research aims to utilize LLMs as a tool integrated in an end-to-end pipeline, and only a very small effort has been dedicated to studying the capabilities of such models in the context of SQL generation alone. It has been shown that language models trained to follow chain of thought (CoT) steps excel at solving problems that involve math and symbolic reasoning (Mirzadeh et al., 2024; Sprague et al., 2024; Ye et al., 2024a, b), but these findings have not been validated as of yet in the context of the text-to-SQL task. Recent work in the domain of LLM theory paves the way for the discovery of LLM universal laws, pertaining to learning language structures (Allen-Zhu and Li, 2024a), knowledge storage and manipulation (Allen-Zhu and Li, 2024b, c, d), learning from mistakes and the ability to self-correct (Ye et al., 2024a, b). We aim to validate the applicability of the latter to coding LLMs and the text-to-SQL downstream task.
3. Methodology
@@
In this section, we describe our RetrySQL training paradigm in detail. We augment the BIRD benchmark dataset (Li et al., 2023) (Section 3.1) with synthetically generated reasoning steps (Section 3.2). Then, we define the retry data generation process, in which reasoning steps are corrupted with random errors and then corrected (Section 3.3).
3.1. Training data
@@
In order to acquire a sizable training dataset for the text-to-SQL task, we utilized the existing training data from the BIRD benchmark (Li et al., 2023). The dataset includes 9428 examples, each consisting of the database name, natural language question, external evidence, and the ground truth SQL query (SQLite dialect). In addition, the metadata for each database is available as well, consisting of a full list of tables, columns and table relations. We discovered that relations for one table, mondial_geo, are defined incorrectly. We excluded it from our pipeline, which left us with 9135 training examples.
@@
We used DDL for schema representation because it provides a concise notation that includes table and column names, together with data types and relations. Moreover, it keeps an SQL-focused context for the model, without needing to explain specific data formats in the prompt. This approach is commonly used in existing text-to-SQL pipelines (Shi et al., 2024).
3.2. Reasoning step generation
Previous work showed that the usage of retry data in model pre-training is effective only if we also instruct the model to follow a chain of reasoning steps (Ye et al., 2024b). To this end, we needed to procure reasoning steps for each of our training examples. The BIRD training dataset does not contain this data, so we used GPT-4o for generating synthetic reasoning steps (Fig. 1a). Enhancing language model training data with synthetically generated components is a newly emerging trend (Abdin et al., 2024). We used a prompt that highlighted the need of reasoning steps being in a format resembling solution reasoning chains from the dataset used in previous research on self-correction (Ye et al., 2024a) (Fig. S3). We verified the correctness of the output formatting and also the semantic validity for a small subset of all examples, consisting of 100 instances sampled uniformly at random, thus representing a wide range of databases and query difficulties. We found that there were no cases with erroneous reasoning steps. Consequently, we assumed that the full dataset is similarly error-free. Full manual verification was not necessary, since we did not aim for the reasoning steps to be perfectly accurate. We ultimately wanted to generate SQL queries, and the reasoning steps were meant to serve as an additional training signal to the verified ground truth from the BIRD dataset.
3.3. Retry data generation
We generated SQL-specific retry data by corrupting the solution steps prepared beforehand (following (Ye et al., 2024b)) (Fig. 1b). We considered several variants of perturbations: forward single (denoted as FS), forward and back single (FBS), forward multiple (FM), forward and back multiple (FBM). Given a sequence of reasoning steps of length $N$, for each step $r_{i}$ in that sequence we select uniformly at random (with probability $p_{retry}$) another step $r_{error}\in S$, where $S$ is a set of candidate corruptions. The selection is done either once (for FS and FBS) or multiple times (for FM and FBM). For FS and FM, $S$ consists of elements $r_{i+1},...\ ,r_{N}$ (i.e. future steps). For FBS and FBM, $S$ contains elements $r_{1},...\ ,r_{i-1},r_{i+1},...\ ,r_{N}$ (i.e. future and past steps). After selecting an element $r_{error}$ from $S$, we follow it with the [BACK] token and then with the correct $r_{i}$ itself. As such, each $r_{i}$ can be replaced with the following:

### v2

@@ Abstract
The text-to-SQL task is an active challenge in Natural Language Processing. Many existing solutions focus on using black-box language models extended with specialized components within customized end-to-end text-to-SQL pipelines. While these solutions use both closed-source proprietary language models and coding-oriented open-source models, there is a lack of research regarding SQL-specific small generative models. At the same time, recent advancements in self-correcting generation strategies show promise for improving the capabilities of existing architectures. The application of these concepts to the text-to-SQL task remains unexplored. In this paper, we introduce RetrySQL, a new approach to training text-to-SQL generation models. We prepare reasoning steps for reference SQL queries and then corrupt them to create retry data that contains both incorrect and corrected steps, divided with a special token. We continuously pre-train open-source coding models with this data and demonstrate that retry steps yield an improvements of up to 4 and 9 percentage points for overall and challenging execution metrics, respectively, as compared to pre-training without retry data. We showcase that the self-correcting behavior is learned by the model and the increase in downstream accuracy metrics is a result of this additional skill. Finally, we incorporate RetrySQL-trained models into the full text-to-SQL pipeline and showcase that they are competitive in terms of Execution Accuracy with proprietary models that contain orders of magnitude more parameters. RetrySQL demonstrates that self-correction can be learned in the text-to-SQL task and provides a novel way of improving generation accuracy for small SQL-oriented language models.
Codehttps://github.com/allegro/RetrySQL
Extended versionhttps://arxiv.org/abs/2507.02529
1 Introduction
@@
The task of translating natural language questions to SQL queries is a major challenge for machine learning models. The complexity stems from the need of relating often ambiguous user input to abstract entities, relations and values that are present in relational databases (li_can_2023). Even prominent Large Language Models (LLMs), such as GPT-4o (openai_2024) or Gemini 1.5 (gemini_2024), struggle with approaching human performance in leading text-to-SQL benchmarks: BIRD (li_can_2023) and SPIDER 2.0 (lei_spider2_2024). The same is true for models tuned specifically for coding tasks (li_codes_2024; chen2024_opensql; talaei_chess_2024). Thus, there exists a need for more advanced solutions that can bridge that gap and provide reliable SQL queries even in difficult real-world scenarios.
@@
The text-to-SQL task can be divided into three main steps (maamari_death_2024): retrieval, generation and correction. Many existing approaches try to tackle these steps at the same time, in a single end-to-end pipeline (maamari_death_2024; talaei_chess_2024; pourreza_din-sql_2023; pourreza_dts-sql_2024), utilizing relatively large LLMs. In this work, we focus only on the generation step and show how it can be improved with a novel approach to model pre-training of small LLMs - text-to-SQL systems need to operate quickly in real-world scenarios.
@@
Figure 1: RetrySQL overview. (a) Reasoning step generation. For each SQL query in the training dataset, we generate a series of reasoning steps using GPT-4o. (b) Preparation of retry data. For each set of reasoning steps, we apply random perturbations, treated as errors, by replacing some steps with different ones. We follow these errors with special [BACK] tokens and amend them with correct steps. (c) We take an open-source LLM and continue its pre-training with training examples that contain retry data injected into reasoning steps. The resulting RetrySQL-trained model learns the ability to self-correct, which improves its capabilities in generating correct SQL queries from natural language questions.
@@
Specifically, we teach the model to self-correct during the generation itself. While previous work did use self-correction in the sense of post-processing, applied in the correction step (pourreza_din-sql_2023), we enforce the self-correcting behavior at an earlier point. This sort of active knowledge-based self-correction is an ongoing research area when it comes to LLMs (ye_physics_part2point1_2024; ye_physics_part2point2_2024). While recent work in slow thinking reasoning systems, such as DeepSeek-R1 (deepseekai_2025), shows that self-correction can be learned in a reinforcement learning setup, other lines of research suggest that it is possible to obtain the self-correction ability with specific data augmentations and a standard auto-regressive pre-training objective (ye_physics_part2point1_2024; ye_physics_part2point2_2024) or during supervised fine-tuning (muennighoff_s1_2025). It has been shown that augmenting training data for grade-school math solution generation with so-called retry data leads to increased generation accuracy (ye_physics_part2point2_2024). The applicability of this approach to other tasks and models has not been explored as of yet.
@@
We introduce RetrySQL, a novel text-to-SQL generation module training paradigm that incorporates retry data in the training process and teaches the resulting model to self-correct. RetrySQL first augments the training data with reasoning steps that explain the sequence of operations required for obtaining the solution SQL query (Fig. 1a). Then, retry data is generated by corrupting the order of these reasoning steps (Fig. 1b). The retry data is incorporated into the training examples and we perform continued pre-training of open-source coding-oriented LLMs, which results in a RetrySQL-trained models that are capable of self-correction (Fig. 1c). We present a set of experiments across multiple strategies of generating reasoning corruptions. We demonstrate that using retry data yields superior generation results when compared to training data with error-free reasoning steps.
@@
To showcase that the self-correcting behavior is indeed learned by the model, we provide an analysis of output token confidence. We show that the max softmax score is on average lower for tokens before correction than for those after. Similarly, incorrect tokens display higher variance of softmax scores across beam search passes than the corrected ones. This illustrates that the model becomes uncertain as it makes a mistake and then self-corrects itself with higher confidence.
@@
Our results corroborate the recent findings regarding the self-correction ability in language models (ye_physics_part2point2_2024), demonstrating that the improvement in generation accuracy coming from the inclusion of retry data in pre-training is a universal law. It is applicable not only to the grade school math reasoning problem and GPT-2, but also to the text-to-SQL domain and larger, more modern Transformer-based decoder-only models. These findings suggest that retry data could be adapted to even more domains, especially if reasoning steps can be added to the training examples.
@@
While the main focus of our work is on the SQL generation step in isolation, we also showcase that relatively small, 1.5B-parameter open-source coding models trained with our RetrySQL paradigm are competitive with much larger closed-source proprietary LLMs when used as a part of the full text-to-SQL pipeline. We share all code for retry data generation, as well as model training and evaluation.
@@
  • We show that using retry data in pre-training is beneficial to the generation process, as indicated by the Execution Accuracy metric calculated for the BIRD and SPIDER benchmark datasets.
@@
  • We illustrate that within a simple end-to-end text-to-SQL pipeline, RetrySQL-trained 1.5B-parameter open-source coding models are competitive with proprietary models such as GPT-4o-mini and GPT-4o, which opens new possibilities for employing such small models in real-world text-to-SQL systems.
2 Related work
@@
Early text-to-SQL methods relied on sequence-to-sequence frameworks, using models like Graph Neural Networks, Recurrent Neural Networks, and pre-trained Transformers for encoding queries and schemas (cai_sadga_2021; cao_lgesql_2021; wikisql_2019), while employing slot-filling or auto-regressive decoding to generate SQL queries (ryansql_2020; wang_rat-sql_2020). Recently, the field has shifted with the emergence of LLMs, which are currently leading in the most popular benchmarks (shi_survey_2024). While initial efforts were focused on optimizing prompt designs that leveraged in-context learning (nan2023enhancingfewshottexttosqlcapabilities; gao2023_dailsql) and multi-stage prompting (pourreza_din-sql_2023), the current state-of-the-art is represented by LLM-based pipelines. These latest approaches integrate LLMs in more complex sequences of processing stages, with separate components for schema linking, self-correction, self-debugging, and self-consistency (talaei_chess_2024; pourreza2024_chasesql; lee2024_mscsql; maamari_death_2024; sun2024_sqlpalm).
@@
Compared to closed-source model prompting approaches, open-source model fine-tuning for the text-to-SQL task remains relatively unexplored (shi_survey_2024). Many of the existing works favor parameter-efficient fine-tuning over full-parameter fine-tuning due to the former’s superior training efficiency and lower training costs (shi_survey_2024; chen2024_opensql; zhang2024_sqlfuse; zhang2024_finsql). While the majority of practitioners choose to use powerful general-purpose LLMs as their base models (pourreza_dts-sql_2024; xie2024_deasql; shi_survey_2024), promising results have also been shown by adapting coding LLMs to the text-to-SQL domain (li_codes_2024; chen2024_opensql; talaei_chess_2024), demonstrating that starting from a model already heavily pre-trained on coding tasks, with SQL-related training data, leads to higher performance in benchmark evaluations.
The usage of small LLMs for the text-to-SQL task is an important practical concern, as response time and costs are crucial factors in real-world applications. Existing work explored models as small as 7B parameters (pourreza_dts-sql_2024), but the utilization of even smaller LLMs remains relatively unexplored.
It has been shown that language models trained to follow chain of thought (CoT) (wei_chain_2023) steps excel at solving problems that involve math and symbolic reasoning (GSM; sprague2024_tocot_or_nottocot; ye_physics_part2point1_2024; ye_physics_part2point2_2024; muennighoff_s1_2025). Recent work in the domain of LLM theory paves the way for the discovery of LLM universal laws, including the ability to self-correct (ye_physics_part2point1_2024; ye_physics_part2point2_2024). Test-time compute methods are an active area of research as well, with some methods indicating that reinforcement learning approaches might not be needed for eliciting self-reflection in LLMs (madaan_self-refine_2023; muennighoff_s1_2025). When it comes to small reasoning models, it has been shown that data curation (gunasekar_textbooks_2023) and knowledge distillation (mitra_orca2_2023) are effective techniques for improving model performance.
The intersection of the text-to-SQL task and reasoning with self-correction for small ($\sim$1B parameters) LLMs has not been explored as of yet. RetrySQL addresses this gap by providing a method for teaching such small text-to-SQL models to self-correct, which greatly improves their capabilities.
3 Methodology
@@
In this section, we describe our RetrySQL training paradigm in detail. We augment the training dataset (Section 3.1) with synthetically generated reasoning steps (Section 3.2). Then, we define the retry data generation process, in which reasoning steps are corrupted with random errors and then corrected (Section 3.3).
3.1 Training data
@@
In order to assess the efficacy of RetrySQL, we utilized two popular text-to-SQL benchmark datasets: BIRD (li_can_2023) and SPIDER (yu_spider_2019). We selected them due to their relatively large training sets. For BIRD, the training data includes 9428 examples, each consisting of the database name, natural language question, external evidence, and the ground truth SQL query (SQLite dialect). In addition, the metadata for each database is available as well, consisting of a full list of tables, columns and table relations. We discovered that relations for one table, mondial_geo, are defined incorrectly. We excluded it from our pipeline, which left us with 9135 training examples. For SPIDER, the training data consists of 8659 examples, with a similar structure to BIRD, excluding the external evidence. We did not use the newer SPIDER 2.0 (lei_spider2_2024) dataset due to its focus on agentic evaluation - it does not include any training examples.
@@
We used DDL for schema representation because it provides a concise notation that includes table and column names, together with data types and relations. Moreover, it keeps an SQL-focused context for the model, without needing to explain specific data formats in the prompt. This approach is commonly used in existing text-to-SQL pipelines (shi_survey_2024).
3.2 Reasoning step generation
Previous work showed that the usage of retry data in model pre-training is effective only if we also instruct the model to follow a chain of reasoning steps (ye_physics_part2point2_2024). To this end, we needed to procure reasoning steps for each of our training examples. The training datasets do not contain this data, so we used GPT-4o for generating synthetic reasoning steps (Fig. 1a). Enhancing language model training data with synthetically generated components is a newly emerging trend (abdin_phi_2024). We used a prompt that highlighted the need of reasoning steps being in a format resembling solution reasoning chains from the dataset used in previous research on self-correction (ye_physics_part2point1_2024) (Fig. S3). We verified the correctness of the output formatting and also the semantic validity for a small subset of all examples, consisting of 100 instances sampled uniformly at random, thus representing a wide range of databases and query difficulties. We found that there were no cases with erroneous reasoning steps. Consequently, we assumed that the full dataset is similarly error-free. Full manual verification was not necessary, since we did not aim for the reasoning steps to be perfectly accurate. We ultimately wanted to generate SQL queries, and the reasoning steps were meant to serve as an additional training signal to the verified ground truth from the training datasets.
3.3 Retry data generation
We generated SQL-specific retry data by corrupting the solution steps prepared beforehand (ye_physics_part2point2_2024) (Fig. 1b). We considered several variants of perturbations: forward single (denoted as FS), forward and back single (FBS), forward multiple (FM), forward and back multiple (FBM). Given a sequence of reasoning steps of length $N$, for each step $r_{i}$ in that sequence we select uniformly at random (with probability $p_{retry}$) another step $r_{error}\in S$, where $S$ is a set of candidate corruptions. The selection is done either once (for FS and FBS) or multiple times (for FM and FBM). For FS and FM, $S$ consists of elements $r_{i+1},...\ ,r_{N}$ (i.e. future steps). For FBS and FBM, $S$ contains elements $r_{1},...\ ,r_{i-1},r_{i+1},...\ ,r_{N}$ (i.e. future and past steps). After selecting an element $r_{error}$ from $S$, we follow it with the [BACK] token and then with the correct $r_{i}$ itself. As such, each $r_{i}$ can be replaced with the following: